OpenStudy (anonymous):

a. Use implicit differentiation to find the derivative dy/dx b. Find the slope of the curve at the given point y^4=6x ; (216,6)

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

help with b, please

OpenStudy (jennychan12):

use point slope formula plug in x and y into the implicitly differentiated equation and then use point slope formula to find the equation

OpenStudy (anonymous):

\[\frac{ 3 }{ 2y ^{3} }\]

OpenStudy (jennychan12):

plug in your x and y values into \[\frac{ dy }{ dx} = \frac{ 3 }{ 2y^3 }\]

OpenStudy (jennychan12):

jkjk just your y value

OpenStudy (anonymous):

1/24?

OpenStudy (jennychan12):

check your math :0

OpenStudy (anonymous):

1/48?

OpenStudy (jennychan12):

your derivative is 3/2y^3

OpenStudy (anonymous):

o duh, im doing it by the ^2. sooo 1/576?

OpenStudy (jennychan12):

lol yes. uhhh no a little bigger than that.

OpenStudy (anonymous):

how is it so?

OpenStudy (jennychan12):

it's 1/144...?

OpenStudy (anonymous):

im confused. 3/(2(6))^3

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):